The system of equations is coincident.

What are the missing values?
3x+y=7
9x+__y=__

Answer by TimothyLamb(4379)

3x + y = 7
3(3x + y = 7)
9x + 3y = 21
---
the missing values are:
3
21
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---